# Population Transformer: Learning Population-level Representations of Intracranial Activity

### v1

@@ 3 Population Transformer Approach
Table 1: Pretraining PopT is critical to downstream decoding performance. We test on a variety of audio-linguistic decoding tasks (see Section 4) with either a single channel (row 1) or 90 channels (rows 2-5) as input. Shown are the ROC-AUC mean and standard error across subjects. We see that all aggregation approaches (rows 2-5) outperform single-channel decoding with BrainBERT [8] (row 1). Pretraining PopT and then fine-tuning it for downstream decoding results in significantly better performance (bold) compared to non-pretrained aggregation approaches (rows 2-4). This gain cannot be explained by simply providing more temporal embeddings, as evidenced by the performance of Linear and Deep NN (rows 2 and 3) that take the concatenated raw temporal embeddings as input. Neither can the gain be attributed to simply using a Transformer architecture, as is shown by a comparison with a non-pretrained PopT (row 4).
@@
Figure 2: Pretrained PopT downstream performance scales better with ensemble size. Increasing channel ensemble size from 1 to 30 (x-axis), we see pretrained PopT (green) decoding performance (y-axis) not only beat non-pretrained approaches (orange, purple, pink), but also continually improve more with increasing channel count. Shaded bands show the standard error across subjects.
Pretraining Our pretraining objective has two components: channel-wise discrimination and next brain state discrimination, which is a group-level objective (see Figure 1b). First, we describe the next brain state discrimination task. Two different subsets of channels $S_{A},S_{B}\subset C$ are chosen with the condition that they be disjoint $S_{A}\cap S_{B}=\emptyset$. During pretraining, the model receives the activities from these channels at separate times $X_{A}=\{x^{t}_{i}\mid i\in S_{A}\}$ and $X_{B}=\{x^{t^{\prime}}_{i}\mid i\in S_{B}\}$. The objective of the task is then to determine whether these states $X_{A}$ and $X_{B}$ have occurred consecutively in time or are separated by some further, randomly selected interval. Given the output of the classification head, the objective is the binary cross entropy: $\frac{1}{N_{\text{batch}}}\sum_{i}y^{*}_{i}\mathcal{L}_{N}=y^{*}_{cls}\log(p(\tilde{y}_{cls}))+(1-y^{*}_{cls})\log(p(\tilde{y}_{cls}))$ where $y^{*}_{cls}=\mathbf{1}(|t-t^{\prime}|<500ms)$.
Next we describe our channel-wise discriminative learning. The token level objective is to determine whether a channels activity has been swapped with activity from a random time. Precisely, activity from each channel $i$ is drawn from a time $t_{i}$. All channels are drawn from the same time $t_{i}=T$, and then 10% of the channels are randomly selected to have their activity replaced with activity from the same channel, but taken from a random point in time $t_{i}\neq T$. Then, given the token outputs of the Population Transformer, the objective function is the binary cross entropy: $\mathcal{L}_{C}=\frac{1}{N_{\text{batch}}}\sum_{i}y^{*}_{i}\log(p(\tilde{y}))+(1-y^{*}_{i})\log(p(\tilde{y}_{i}))$ where $y^{*}_{i}=\mathbf{1}(t_{i}\neq t)$.
Then, our complete objective function is $\mathcal{L}=\mathcal{L}_{N}+\mathcal{L}_{C}$.
Fine-tuning During fine-tuning, the [CLS] intermediate representation, $\tilde{y}_{cls}$ of the pretrained PopT is passed through a single layer linear neural network to produce a scalar $\hat{y}_{cls}$. This scalar is the input to binary cross entropy loss for our decoding tasks (see Section 4).
4 Experiment Setup
Data We use the publicly available subject data from [8]. Data was collected from 10 subjects (total 1,688 electrodes, with a mean of 167 electrodes per subject) who watched 26 movies while intracranial probes recorded their brain activity. The movie transcripts were aligned to the brain activity so that features such as volume, pitch, etc. could be associated with the corresponding sEEG readings. 19 of the sessions are used for pretraining. 7 of the sessions are held-out for evaluation.
Decoding We evaluate the effectiveness of our pretrained PopT model by fine-tuning it on the four downstream decoding task used in the evaluation of [8]. Two of the tasks are audio focused: determining whether a word is spoken with a high or low pitch and determining whether a word is spoken loudly or softly. And two of the tasks have a more linguistic focus: determining whether the beginning of a sentence is occurring or determining whether any speech at all is occurring.
Our approach enables decoding on any arbitrary size of ensemble. We verify that our model is able to leverage additional channels for improved decoding performance that scales the number of inputs. To test this, we first order the electrodes by their individual linear decodability per task, and we increase the number of channels available to the model at fine-tuning time.
Baselines We want to determine whether the information about spatial relationships learned during pretraining was useful at fine-tuning time. For comparison, we concatenate the raw BrainBERT embeddings and train a linear and deep NN on the decoding task. This sets a baseline for how much improvement is achievable from simply having more channels available at once. To determine whether our performance can be attributed to using a more powerful architecture, we also fine-tune a PopT without pretraining, i.e. with randomly initialized weights.
@@ 5 Results
Decoding performance Compared to trying to decode from the bare BrainBERT embeddings or from a non-pretrained PopT, the PopT both achieves better decoding performance (see Table 1) and does so with steeper scaling per added channel (Figure 2).
@@
To verify that the weights of the pretrained PopT capture neural processing well even without fine-tuning, we also train a linear-encoder on top of the frozen PopT [CLS] token and find the same trends ( Figure 10: Frozen scaling – Figure 10). This point in particular is important in building confidence in the results of our interpretability studies (see Section 6), in which we use the frozen pretrained weights to analyze connectivity.
@@
Sample and compute efficiency Our PopT learns spatial relationships between channels, in a way that makes downstream supervised learning more data and compute efficient (see Figure 3 and Figure 4). Compared to the non-pretrained baseline models, fine-tuning the pretrained PopT achieves more decoding performance from fewer samples. At only 200 examples, the pretrained PopT has already surpassed the performance achieved by the non-pretrained model on the full dataset, for the volume, sentence onset, and speech vs. non-speech tasks Figure 3. The number of steps required for each model to converge is also greatly reduced by starting with the pretrained PopT. We see that fine-tuning the pretrained PopT consistently requires 500 steps or fewer steps to reach its converged performance Figure 4, whereas the non pretrained baselines may require 2k or more steps.
@@
Figure 3: Pretrained PopT is more sample efficient when fine-tuning. Varying the number of samples available to each model at train time (x-axis), we see how the pretrained PopT is highly sample efficient, requiring only a fraction of samples to reach the full performance level of non pretrained aggregation approaches (dashed lines). Bands show standard error across test subjects. Stars indicate performance of the model trained on the full fine-tuning dataset.
@@
Figure 4: Pretrained PopT is consistently more compute efficient when fine-tuning. Number of steps required for each model to reach final performance during fine-tuning (dashed lines). We find that pretrained PopT consistently requires fewer than 750 steps (each step is training on a batch size of 256) to converge, in contrast to the 2k steps required for the non pretrained PopT. Linear aggregation can be similarily compute efficient, but occasionally benefits from more training steps depending on dataset size (Speech vs. Non-speech). Bands show standard error across test subjects. Stars indicate fully trained performance.
@@
Figure 5: Gains in decoding performance are available to new subjects. To test whether our pretrained PopT weights will be able to yield decoding benefits for unseen subjects, we run a hold-one-out analysis in which we exclude one subject from pretraining and then evaluate on that subject during fine-tuning (Held-out). We compare this with our full PopT model that has seen all subjects during pretraining (All). A minimal decrease in downstream decoding performance is found if the subject is held-out from pretraining (Held-out vs All). This is in stark contrast to the achievable downstream performance with a non pretrained PopT (Non-pretrained PopT).
@@
Generalizability To test whether our pretrained weights will be useful for subjects not seen during training, we conduct a hold-one-out analysis. We pretrain a model using all subjects except for one, and then fine-tune and evaluate on the model downstream. We find that missing a subject from pretraining does not significantly affect the downstream results (see Figure 5). This raises our confidence that the pretrained weights will be useful for unseen subjects and for researchers using new data.
@@
Ablation of loss components and position information An ablation study confirms that both the network-wise and channel-wise component of the pretraining objective contribute to the downstream performance (Table 2). We also find that including the 3D position information for each channel is critical for decoding. These findings also hold when the PopT is kept frozen during fine-tuning (see Appendix G: Frozen ablation – Table 4). Additionally, we find that the discriminative nature of our loss is necessary for decoding. Attempting to add an L1 reconstruction term to our pretraining objective results in poorer performance, perhaps because the model learns to overfit on low-entropy features in the embedding. Our discriminative loss requires the model to understand the embeddings in terms of how they can be distinguished from one another, which leads the model to extract more informative representations.
Figure 6: Pretraining with more subjects leads to better downstream performance. We pretrain PopT with different number of subjects (colors) and test on our decoding tasks (x-axis). Bars indicate mean and standard error of performance across channel ensembles 5-30 on test subject 3. Model descriptions: 0 subjects (non-pretrained), 1 subject (pretrain w/ subject 4), 2 subjects (pretrain w/ subjects 4, 8), 3 subjects (pretrain w/ subjects 4, 8, 10), All subjects (pretrain w/ all 10 subjects). Pretraining with one subject gives a considerable benefit compared to no pretraining (red to yellow), but the addition of more subjects to pretraining consistently improves performance (yellow $\rightarrow$ green).
@@
Table 2: PopT ablation study. We individually ablate our losses and positional encodings during pretraining then decode on the resulting models. Shown are ROC-AUC mean and standard error across subjects. The best performing model across all decoding tasks uses all three of our proposed components, showing that they are all necessary. Removing our positional encoding during pretraining and fine-tuning drops the performance the most, indicating that position encoding is highly important for achieving good decoding. Additionally, we attempt adding a reconstruction component to the loss as a regularizing term, but find that this leads to poorer performance (see section 5).
Figure 7: Probing the pretrained model for inter-channel connectivity Traditionally, connectivity analysis between regions is done by computing the coherence [37], i.e. cross-correlation, between electrode activity (left). We propose an alternative analysis based on how channels matter to each other in the context of our pretraining objective. Iteratively, we select an electrode, mask out its activity, and then plot the degradation in the channel-wise objective function of the pretrained PopT objective for the unmasked electrodes. Plotting the values of this delta (right) recovers the main points of connectivity, purely based off of the relationships learned during pretraining. Shown here is a plot for a single subject; plots for all test subjects can be seen in Appendix E: Connectivity.
@@ 6 Interpreting Learned Weights
Our final analysis are two interpretability studies of the Population Transformer’s learned weights. In the first, we use the PopT weights to uncover connectivity maps of the channels, and in the second, we use the attention weights of the fine-tuned PopT to identify candidate functional brain regions per decoding task.
Connectivity For identifying connectivity per region, traditional neuroscience analyses typically use cross-correlation as a measure of channel connectivity [37]. Our PopT allows for an alternative method of determining connectivity, based on the degree to which channels are sensitive to each other’s context. We go through our channels, masking one channel and then evaluating the model’s performance on the pretraining channel-wise objective for the remaining unmasked channels. We take the degradation in performance as a measure of connectivity. We can construct plots as in Figure 7, that recapitulate the strongest connectivity of the cross-correlation maps. Note that while some approaches for modelling brain activity explicitly build this into their architecture [25] we recover these connections purely as a result of our self-supervised learning.
Candidate functional brain regions from attention weights Next, we discuss the possibility of uncovering functional brain regions from the attention weights. After fine-tuning our weights on a decoding task, we can examine the attention weights of the [CLS] output for candidate functional brain regions. We obtain a normalized Scaled Attention Weight metric across all subjects to be able to analyze candidate functional brain regions across sparsely sampled subject datasets Figure 8. The Scaled Attention Weight is computed from raw attention weights at the [CLS] token passed through the attention rollout algorithm [38]. The resulting weights from each channel are then grouped by brain region according to the Destrieux layout [39]. Additional details available in Appendix D.
The resulting weights reveal expected functional brain regions related to the tasks decoded Figure 8. For our low-level perceptual auditory tasks (Volume and Pitch), we see that our model learns to attend to the primary auditory cortex. For our higher-level language distinction tasks (Speech vs. Non-speech and Sentence onset), we see higher attention is placed at language areas like Wernicke’s area. Given the massive pretraining PopT undergoes, these scaled attention weights provide a valuable a new tool for discovering candidate functional brain regions.
Figure 8: Attention weights from a fine-tuned PopT identify candidate functional brain regions Candidate functional maps can be read from attention weights of a PopT fine-tuned on our decoding tasks. For the Volume and Pitch tasks, note the weight placed on the primary auditory cortex (black arrows), but not in Wernicke’s area. For the Speech vs Non-speech and Sentence onset tasks, note the weight placed on regions near Wernicke’s area (black arrows). Center brain figure highlight regions related to auditory-linguistic processing such as language production area Broca’s area, language understanding Wernicke’s area, and the primary auditory cortex (adapted from [40]).

### v2

@@ 4 Experiment Setup
TOTEM:
      Linear Agg. $0.55\pm 0.02$ $0.66\pm 0.03$ $0.79\pm 0.04$ $0.77\pm 0.05$
      Deep NN Agg. $0.57\pm 0.02$ $0.67\pm 0.03$ $0.78\pm 0.03$ $0.75\pm 0.05$
      Non-pretrained PopT $0.53\pm 0.02$ $0.64\pm 0.02$ $0.79\pm 0.03$ $0.77\pm 0.05$
      Pretrained PopT $\mathbf{0.60}\pm\mathbf{0.02}$ $\mathbf{0.73}\pm\mathbf{0.02}$ $\mathbf{0.86}\pm\mathbf{0.03}$ $\mathbf{0.84}\pm\mathbf{0.06}$
End-to-end:
      Brant (Zhang et al., 2024) $\mathbf{0.61}\pm\mathbf{0.03}$ $\mathbf{0.74}\pm\mathbf{0.03}$ $\mathbf{0.80}\pm\mathbf{0.04}$ $\mathbf{0.80}\pm\mathbf{0.03}$
Table 1: Pretraining PopT is critical to downstream decoding performance (iEEG data). We test on a variety of audio-linguistic decoding tasks (see Section 4) with 90 channels as input. The temporal encoder used for aggregation in sections 1 and 2 are denoted in the section header. We also evaluate against an end-to-end pretrained iEEG model in section 3. Shown are the ROC-AUC mean and standard error across subjects. Best per section are bolded.
Model Balanced Accuracy ROC AUC
Chronos
      Linear Agg. $0.7754\pm 0.0008$ $0.8563\pm 0.0003$
      Deep NN Agg. $0.7881\pm 0.0057$ $0.8678\pm 0.0049$
      Non-pretrained PopT $0.7763\pm 0.0047$ $0.8631\pm 0.0016$
      Pretrained PopT $\mathbf{0.7976}\pm\mathbf{0.0022}$ $\mathbf{0.8821}\pm\mathbf{0.0016}$
TS2Vec
      Linear Agg. $0.7649\pm 0.0005$ $0.8533\pm 0.0003$
      Deep NN Agg. $0.7853\pm 0.0021$ $0.8721\pm 0.0015$
      Non-pretrained PopT $0.7896\pm 0.0037$ $0.8782\pm 0.0018$
      Pretrained PopT $\mathbf{0.8063}\pm\mathbf{0.0010}$ $\mathbf{0.8907}\pm\mathbf{0.0019}$
End-to-end:
      BIOT (Yang et al., 2024) $0.7959\pm 0.0057$ $0.8815\pm 0.0043$
      LaBraM (Jiang et al., 2024) $\mathbf{0.8258}\pm\mathbf{0.0011}$ $\mathbf{0.9162}\pm\mathbf{0.0016}$
Table 2: Pretraining PopT is critical to downstream decoding performance (EEG data). We test on a seizure detection task (TUAB in Obeid & Picone (2016)) with 21 channels as input. The temporal encoder used for aggregation in sections 1 and 2 are denoted in the section header. We also evaluate against end-to-end pretrained EEG models in section 3 (values from the original works). Shown are the ROC-AUC mean and stdev across 5 random seeds. Best per section are bolded.
Figure 2: Compared to common aggregation approaches, pretrained PopT consistently yields better downstream decoding across tasks, data modalities, and temporal embedding types. NPopT = Non-pretrained PopT. (a) performance on four audio-linguistic iEEG tasks with 90 electrodes. Grey bars denote standard error across subjects. (b) performance on a seizure detection EEG task with 21 electrodes. Grey bars denote standard deviation across 5 random seeds.
@@
Figure 3: Pretrained PopT downstream performance scales better with ensemble size. Increasing channel ensemble size from 1 to 50 (x-axis), we see pretrained PopT (green) decoding performance (y-axis) not only beat non-pretrained approaches (orange, purple, grey), but also continually improve more with increasing channel count. Shaded bands show the standard error across subjects.
@@ 5 Results
Decoding performance We find that using a pretrained PopT significantly benefits downstream decoding compared to baseline channel aggregation techniques across tasks, data modalities, and temporal encoding models (Tables 1, 2 and 2). The pretrained PopT scales well with increasing ensemble sizes (Figure 3a), a challenging task for the baseline aggregation approaches due to limited downstream task data and increasing input size. To test our method’s ability to handle multiple types of channel encodings, we applied our framework to 4 different channel encoders: (1) an iEEG-specific temporal encoder (BrainBERT (Wang et al., 2022)), (2) a general tokenization-based time-series encoder (TOTEM (Talukder et al., 2024)), (3) a pretrained general time-series encoder (Chronos (Ansari et al., 2024)), and a general convolution-based time-series encoder (TS2Vec (Yue et al., 2022)). We see significant improvements in performance with the pretrained PopT in all cases when comparing with baseline aggregation approaches (Figure 2).
We also find that PopT can achieve competitive performance against pretrained end-to-end models, such as Brant (Zhang et al., 2024) for iEEG, and BIOT (Yang et al., 2024) and LaBraM (Jiang et al., 2024) for EEG (Tables 1 and 2). For instance, PopT outperforms Brant (Zhang et al., 2024) in decoding iEEG data with our pretrained PopT + BrainBERT combination, likely due to PopT’s ability to leverage spatial relationships. Whereas Brant leaves the channel aggregation problem open. PopT is competitive with recent end-to-end trained EEG models (Yang et al., 2024; Jiang et al., 2024) at the EEG TUAB seizure detection task. This is impressive, since models such as LaBraM were specifically developed for this application, whereas PopT was trained on top of generic time-series embeddings. We find that PopT can offer an efficient and competitive alternative to large end-to-end models for these decoding tasks, due to the effectiveness of our pretraining task for learning spatial and functional relationships between channel input embeddings.
@@
To verify that the weights of the pretrained PopT capture neural processing well even without fine-tuning, we also train a linear-encoder on top of the frozen PopT [CLS] token and find the same trends (Figure 11. This point in particular is important in building confidence in the results of our interpretability studies (see Section 6), in which we use the frozen pretrained weights to analyze connectivity. Finally, for the remaining analyses described below, we use a PopT with BrainBERT inputs.
@@
Sample and compute efficiency Our PopT learns spatial relationships between channels, in a way that makes downstream supervised learning more data and compute efficient (see Figure 4 and Figure 5). Compared to the non-pretrained baseline models, fine-tuning the pretrained PopT can achieve the same decoding performance as other aggregation techniques with an order of magnitude fewer samples. The pretrained PopT surpasses the performance achieved by all other aggregation techniques by 500 samples out of the full dataset (roughly 5-10k examples depending on subject and task) Figure 4. The pretrained PopT also converges at a low number of steps, competitive with Linear and Deep NN compute efficiency. This greatly contrasts with the non-pretrained PopT and occasionally the Linear and Deep NN baselines, which may require 2k or more steps (Figure 5).
@@
Figure 4: Pretrained PopT is more sample efficient when fine-tuning. Varying the number of samples available to each model at train time (x-axis), we see that the pretrained PopT is highly sample efficient, requiring only a fraction of samples (fewer than 500 samples out of 5-10k of the full dataset) to reach the full performance level of baseline aggregation approaches (dashed lines). Bands show standard error across test subjects. Stars indicate performance with full fine-tuning dataset.
@@
Figure 5: Pretrained PopT is consistently compute efficient when fine-tuning. Number of steps required for each model to reach final performance during fine-tuning (dashed lines). The pretrained PopT consistently requires fewer than 750 steps (each step is an update on a batch size of 256) to converge, in contrast to the 2k steps required for the non-pretrained PopT. Linear and Deep NN aggregation can be similarly compute efficient, but occasionally require more training steps depending on task. Bands show standard error across subjects. Stars indicate fully trained performance.
@@
Figure 6: Gains in decoding performance are available to new subjects. A minimal decrease in downstream decoding performance is found if the subject is held-out from pretraining (Held-out vs All). This is in stark contrast to the achievable downstream performance with a non-pretrained PopT (Non-pretrained PopT).
@@
Generalizability To test if our pretrained weights will be useful for subjects not seen during training, we conduct a hold-one-out analysis. We pretrain a model using all subjects except for one, and then fine-tune and evaluate on the model downstream. We find that missing a subject from pretraining does not significantly affect the downstream results (see Figure 6). This raises our confidence that the pretrained weights will be useful for unseen subjects and for researchers using new data.
@@
Ablation of loss components and position information An ablation study confirms that both the ensemble-wise and channel-wise component of the pretraining objective contribute to the downstream performance (Table 3). Furthermore, including the 3D position information for each channel is critical for decoding. These findings also hold when the PopT is kept frozen during fine-tuning (see Appendix H: Frozen ablation). Additionally, we find that the discriminative nature of our loss is necessary for decoding. Attempting to add an L1 reconstruction term to our pretraining objective results in poorer performance, perhaps because the model learns to overfit on low-entropy features in the embedding. Our discriminative loss requires the model to understand the embeddings in terms of how they can be distinguished from one another, which leads the model to extract more informative representations.
Figure 7: Pretraining with more subjects leads to better downstream performance. We pretrain PopT with different number of subjects (colors) and test on our decoding tasks (x-axis). Bars indicate mean and standard error of performance across channel ensembles 5-30 on a held out test subject. Pretraining with one subject gives a considerable benefit compared to no pretraining (red to yellow), but the addition of more subjects to pretraining consistently improves performance (yellow $\rightarrow$ green).
@@
PopT with L1 reconstruction only $0.56\pm 0.04$ $0.65\pm 0.08$ $0.73\pm 0.10$ $0.74\pm 0.10$
Table 3: PopT ablation study. We individually ablate our losses and positional encodings during pretraining then decode on the resulting models. Shown are ROC-AUC mean and standard error across subjects evaluated at 90 electrodes. The best performing model across all decoding tasks uses all of our proposed components, showing that they are all necessary. Removing our positional encoding during pretraining and fine-tuning drops the performance the most, indicating that position encoding is highly important for achieving good decoding. Additionally, we attempt adding or only using a reconstruction component to the loss as a regularizing term, but find that this leads to poorer performance (last two rows).
Figure 8: Probing the pretrained model for inter-channel connectivity Traditionally, connectivity analysis between regions is done by computing the coherence between electrode activity (left). We propose an alternative analysis purely based on the contextual sensitivity learned during pretraining. Briefly, we select an electrode, mask out its activity, and then measure the degradation in the channel-wise objective function for the remaining electrodes. Plotting the values of this delta (right) recovers the main points of connectivity. Plots for all test subjects can be seen in Appendix F: Connectivity.
@@ 6 Interpreting Learned Weights
Finally, we conduct two interpretability studies of the Population Transformer’s learned weights.
Connectivity Traditional neuroscience analyses typically use cross-correlation as a measure of region connectivity (Wang et al., 2021). Our PopT allows for an alternative method of determining connectivity, based on the degree to which channels are sensitive to each other’s context. In this method, each channel is masked in turn, and then model performance on the pretraining channel-wise objective for the remaining unmasked channels is measured. We use the degradation in performance as a measure of connectivity. We can construct plots as in Figure 8, that recapitulate the strongest connectivity of the cross-correlation maps. Note that while some approaches for modelling brain activity explicitly build this into their architecture (Cai et al., 2023), we recover these connections purely as a result of our self-supervised learning. Additional method details available in Appendix E.
Candidate functional brain regions from attention weights After fine-tuning our weights on a decoding task, we can examine the attention weights of the [CLS] output for candidate functional brain regions. We obtain a normalized Scaled Attention Weight metric across all subjects to analyze candidate functional brain regions across sparsely sampled subject datasets Figure 9. The Scaled Attention Weight is computed from raw attention weights at the [CLS] token passed through the attention rollout algorithm (Abnar & Zuidema, 2020). The resulting weights from each channel are then grouped by brain region according to the Destrieux atlas (Destrieux et al., 2010). A full description of the method is available in Appendix E.
The resulting weights reveal expected functional brain regions related to the tasks decoded Figure 9, with low-level auditory tasks highlighting primary auditory cortex and higher-level language distinction tasks highlighting language-specific areas. Given the massive pretraining PopT undergoes, these scaled attention weights provide a valuable new tool for discovering candidate functional regions.
Figure 9: Attention weights from a fine-tuned PopT identify candidate functional brain regions Candidate functional maps can be read from attention weights of a PopT fine-tuned on our decoding tasks. For the Volume and Pitch tasks, note the weight placed on the primary auditory cortex (black arrows), but not in Wernicke’s area. For the Speech vs Non-speech and Sentence onset tasks, note the weight placed on regions near Wernicke’s area (black arrows). Center brain figure highlight regions related to auditory-linguistic processing (aph, 2017)).